Mr. chittocks political button collection currently has 840 buttons. This represents a 12% increase over his total collection of 3 years ago. Determine how many buttons were in his collection 3 years ago
Answer by ikleyn(52935)

.
Let x be the number three years ago.


According to the condition, this number "x" is the base for the comparison,


so your equation is


    x*(1+0.12) = 840

    1.12x = 840

     x = 840/1.12 = 750.


ANSWER.  3 years ago the collection was 750 buttons.

Solved.

-----------------------

On percentage,  see the lessons
    - Percentage problems
    - Percentage word problems (Type 1 problems, Finding the Part)
    - Percentage word problems (Type 2 problems, Finding the Rate)
    - Percentage word problems (Type 3 problems, Finding the Base)
    - More complicated percentage problems
    - Problems on percentage that lead to unexpected results
in this site.

They are written specially for you.

Read them and become an expert in this area.

Start from the very first lesson of the list.

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.